# Are Large Language Models Rational Investors?

### v1

@@ 3.2 Risk-preference Bias
Asset returns are inherently uncertain, and investor behavior within financial markets is influenced by their risk-return preferences. The exploration of Risk-preference Bias within our framework seeks to elucidate situational dependence bias, loss aversion, and the framing effect across diverse decision-making contexts. It evaluates the consistency of LLMs’ risk preferences amidst varying decision-making scenarios.
@@ 3.2.1 Data Design
In crafting scenarios that mirror real-life decision-making, we assembled a collection of 40 diverse situations spanning career planning, agricultural management, property transactions, and sporting events. Each scenario offers three decision alternatives , where , articulated in terms of potential gains or losses. The presentation order of these alternatives is randomized to minimize option bias in LLMs Zheng et al. (2023).
The construction of these alternatives adheres to the expected utility theory, encapsulated by the equation:
where is the utility function, the outcome under state , and the probability of .
Following the theoretical framework Simon et al. (1994), the utility function’s concavity mirrors risk preferences, with convexity indicating a different disposition. Considering the practical difficulties in direct concavity assessment, we employ the second-order Taylor expansion for utility approximation:
enabling us to modulate risk preferences through variance adjustment. This approach is detailed in Appendix E.
@@ 3.2.2 Analysis Logic
Recognizing decision-making as a process shaped by previous experiences and contextual factors, we delve into the situational dependence bias by examining if LLMs exhibit variable risk preferences across different scenarios . This involves prompting LLMs to choose among gain-framed scenarios and analyzing the pattern of choices across various contexts to discern any situational bias.
Furthermore, to investigate the framing effect, we translate all scenarios from Chinese to English, monitoring shifts in LLM preferences as a function of linguistic framing. This analysis seeks to identify if LLM outputs are influenced by the manner in which scenarios and choices are linguistically constructed, potentially revealing a bias towards the framing of options.
4 Models
In our evaluation, we select BGE333https://huggingface.co/BAAI/bge-large-zh-v1.5 as the Chinese embedding model and test 19 models, including English-oriented, Chinese-oriented, and FinLLMs, such as GPT4 OpenAI (2023), LLaMA Touvron et al. (2023), GLM4 Zeng et al. (2022), Qwen Bai et al. (2023),Xuanyuan Zhang et al. (2023), etc. The detailed settings refer to Appendix D.
5 Results
5.1 Belief Bias
The evaluation of Belief Bias is principally conducted through the examination of event news and interactions. This analysis reveals a widespread Anchoring Effect across the majority of Large Language Models (LLMs) when the subjects of events and interactions are modified, with slight variations observed among different models. The average variance index, detailed in Figure 3, sheds light on the rationality levels of LLMs with respect to representativeness bias. Specifically, LLMs with a focus on the Chinese language, such as the GLM and Qwen series, exhibit commendable financial rationality, whereas the Xuanyuan and Baichuan series are more susceptible to irrational behavior.
In terms of Overconfidence, violin plots, presented in subsection G.1, illustrate the score distributions of various texts across all models. A notable disparity is observed in the models’ responses to composite texts of positive and negative emotional content. As per Table 4, the GPT and InternLM models display a marked optimism, in contrast to the pronounced pessimism of the Qwen and GLM series. Furthermore, Figure 4 highlights that models trained on financial corpora experience a heightened score variability compared to their base counterparts.
In terms of Representativeness Bias, the less rational LLMs demonstrated a correlation coefficient between model scores and market capitalization below 10%, indicating a minimal correlation. However, certain LLMs showed clear biases towards specific industries, as documented in subsection G.1. For example, the FinQwen model consistently allocated lower scores to the Steel and Banking sectors. A comprehensive analysis reveals that the Media, Steel, Banking, and Non-Banking Finance sectors frequently occupy the extreme ends of the scoring spectrum across different models, whereas the Computer Science and Automobile sectors generally maintain a middle ground, exhibiting relative stability.
For the limited Attention Bias, the application of the COT methodology, aimed at slowing down the cognitive process of models, resulted in a minor reduction in rationality among the more logical models but significantly diminished biases in the less rational ones, as shown in subsection G.2. The clustering and keyword analysis indicate that this intentional cognitive strategy improves the logical consistency and financial acumen of the model outputs. A comparison of the reasoning keywords for the top-performing GLM-4 model and the underperforming Baichuan2-7B model, illustrated in Figure 39 and Figure 39 respectively, reveals a more coherent logic in GLM’s reasoning compared to Baichuan2-7B. Moreover, the clustering of the models’ inferential outputs reveals substantial variations in the ratings assigned to each category. This divergence underscores that the irrationality exhibited by models in financial contexts stems more from their inherent cognitive processes than from their computational capabilities.
5.2 Risk-preference Bias
@@
The exploration of Risk-preference Bias entails the examination of LLM decisions across varied scenarios. The compiled results, particularly for gain-framed queries, are tabulated in Table 10. A predominant trend among most models is the exhibition of distinct risk preferences in disparate scenarios, indicative of a pronounced Situational Dependence Bias. Nonetheless, prefacing prompts with an instruction of the model’s risk-averse nature significantly attenuates this bias. In the context of loss-framed queries, some models exhibit a pronounced Loss Aversion Bias like GPT-4, as shown in Table 9. Moreover, refer to Table 8, the translation of all queries into English precipitated notable discrepancies between the models’ responses to Chinese and English versions, underscoring a pronounced Framing Effect.
@@ 6 Discussion
Our study’s contribution through the FBI framework extends to the identification and reduction of irrational tendencies in large-scale models within the finance sector. By employing FBI, researchers can acquire an in-depth comprehension of the biases and logical processes of LLMs in finance. Such insights are pivotal for pinpointing detrimental financial irrationalities and devising strategies to mitigate them. We delve into the influence of Model Size, Training Data, and Input Form modifications on the observed outcomes.
@@ 6.1 Model Size
The findings from subsection 5.1 illustrate a correlation between the enlargement of model parameters and the enhancement of financial rationality, particularly in relation to belief bias. This trend underscores the significance of model size in advancing NLP within financial quantitative analysis. Therefore, our analysis shows that the financial rationality of LLMs often increases with the increase in model size, which is consistent with the scaling law Kaplan et al. (2020).
@@ 6.2 Training Data
The FBI framework evaluated the financial rationality of both general-LLMs and FinLLMs. Insights from Figure 4 indicate that models trained on financial datasets are prone to heightened score variability and an increased inclination towards risk, thereby potentially exacerbating financial irrationality. Meanwhile, based on figures from Figure 12 to Figure 29, we observe that many models exhibit nearly opposite industry biases, such as ChatGLM2-6B and Qwen-7B, yet models from the same period often share similar biases. This suggests that such biases are likely influenced by temporal biases in the training data, in line with the concept of industry cycle rotation. Financial datasets typically encompass research reports, financial statements, corporate disclosures, and financial regulations. Previous studies Wang and Wu (2011) have highlighted the potential for embellishment and over-promotion within online financial content, such as research reports. While incorporating such data can bolster LLM capabilities in NLP tasks specific to finance, it also harbors the risk of embedding financial irrationality. The deployment of these models in financial quantification tasks could inadvertently precipitate financial disturbances, posing a critical concern that necessitates immediate attention.
@@ 6.3 Input Forms
The FBI framework employed three prompt approaches: Direct, Instructional, and COT. Experimental findings from section 5 suggest that the inclusion of instructional prompts or the application of the COT method can enhance the financial rationality of models. Specifically, the Instructional method enables models to adopt predefined risk preference attributes, thereby improving their adherence to directives. Implementing the COT method to emulate a "slow thinking" process effectively moderates the models’ susceptibility to belief bias.
@@ 7 Conclusion
Our research introduces the Financial Bias Indicators (FBI) framework as a pioneering approach to critically assess the financial rationality of Large Language Models (LLMs) within the complex domain of financial analysis. Through a rigorous examination of 19 leading LLMs, our study unveils significant variances in their financial rationality. The findings underscore the pivotal role of model architecture, training paradigms, and input forms in shaping the financial rationality of LLMs. By highlighting the nuances of financial rationality and the impact of inherent biases, this work not only enhances our understanding of LLMs’ capabilities and limitations in financial contexts but also charts a path forward for refining these sophisticated tools. The development of LLMs that are both more financially astute and unbiased is crucial for their reliable application in financial analysis, promising a future where AI-driven insights can be leveraged with greater confidence and precision in the financial sector.
@@ Appendix A Cognitive Bias
| Belief Bias | Limited Attention |
The brain has two systems when working: fast thinking and slow thinking. It uses intuition to deal with things quickly. |
| Situational dependence bias |
The effect of a stimulus depends largely on the context in which it occurs. |
| Loss aversion |
Sensitivity to losses exceeds gains of equal value. |
| Framing effect |
Different descriptions of an objectively identical problem lead to different decision-making judgments. |
| Risk-Preference Bias | Representativeness bias |
When making probability estimates, people tend to focus on certain representative features, ignoring environmental probabilities and sample size. |
| Anchoring effect |
Decision-making is often influenced by the first information received, like an anchor sinking to the bottom of the sea. |
| Overconfidence |
Belief that one’s knowledge is more accurate than the facts; one’s information is given more weight. |
Appendix B Event Type
We have sorted out the types of events that can affect a company’s stock price based on the regular patterns of the Chinese stock market, and finally sorted out four categories, totaling 16 types of events. The detailed content is shown in Table 2.
@@ Appendix B Event Type
| Corporate Governance and Equity Changes |
Major Asset Restructuring |
The process of recombining, adjusting, and allocating the distribution status of enterprise assets among the owners, controllers, and external economic entities. |
Equity Incentive |
By conditionally granting employees partial shareholder rights, a sense of ownership is fostered, forming a community of interests with the company. |
Increase or Decrease in Shareholder Holdings |
Changes in the shareholder holdings of company stocks. |
Buy-back |
The act of a listed company using cash or other means to repurchase its shares from the stock market. |
Circulation of Restricted Stock |
Restricted shares become freely tradable in the secondary market after the commitment period. |
| Financial Reports and Earnings Expectations |
Performance Report |
Regular preparation by each responsibility center to evaluate and assess performance, serving as the basis for future budget preparation. |
| Market Behavior and Announcements |
Private Placement |
Targeted issuance of bonds or stocks to a select group of senior institutional or individual investors. |
Transfer of Shares |
Listed companies transfer their provident fund to share capital in proportion or issue bonus shares accordingly. |
Stock Price Fluctuations |
Sudden large inflows and outflows of funds lead to increased volatility in stock prices. |
Business Dynamics |
Updates on enterprises and their surroundings, using major production and sales information to promote corporate brand and image. |
| Negative Events and Risk Management |
Dispute |
Disputes between companies or between companies and individuals. |
Investigation |
Filing an investigation signifies a basic determination of illegal facts, allowing for compulsory measures and official initiation of investigation procedures. |
Violation Penalties |
Punishments for enterprises violating regulations of regulatory bodies. |
Litigation and Arbitration |
Litigation and arbitration for contract disputes and other property rights disputes between enterprises. |
Security |
Enterprises providing guarantees for loans and other matters for other enterprises. |
Appendix C Company Profile
@@ Appendix D Models
We have selected a total of 19 financial and generic LLMs oriented by Chinese and English, with specific details shown in Table 3.
@@
| Baichuan2-7B | True | 7B | False |
local |
| Baichuan2-13B | True | 13B | False |
local |
| ChatGLM2-6B | True | 6B | False |
local |
| ChatGLM3-6B | True | 6B | False |
local |
| ChatGLM3-Turbo | True | 33B | False |
API |
| GLM-4 | True | Unknown | False |
API |
| InternLM2-7B | True | 7B | False |
local |
| InternLM2-20B | True | 20B | False |
local |
| LLaMA2-7B | False | 7B | False |
local |
| LLaMA2-13B | False | 13B | False |
local |
| Qwen-7B | True | 7B | False |
local |
| Qwen-14B | True | 14B | False |
local |
| FinQwen | True | 14B | True |
local |
| Qwen-72B | True | 72B | False |
local |
| Qwen-max | True | 72B | False |
API |
| Xuanyuan-13B | True | 13B | True |
local |
| Xuanyuan-70B | True | 70B | True |
local |
| GPT-3.5 | False | Unknown | False |
API |
| GPT-4 | False | Unknown | False |
API |

### v2

@@ 5.2.1 Data Design
To simulate real-life decision-making, we designed 40 scenarios with 200 multiple-choice questions, divided into 200 gain-framed and 100 loss-framed scenarios. Each question presents three decision alternatives , where represents risk preferences: Risk-loving, Risk-neutral, Risk-averse. The alternatives are randomized to minimize bias Zheng et al. (2023).
The alternatives are constructed based on expected utility theory Simon et al. (1994), represented by:
where is the utility function, the outcome, and the outcome’s probability.
The concavity of , reflecting risk preferences, is indirectly assessed via second-order Taylor expansion:
This approximation adjusts risk preferences by modulating variance, with technical details in Appendix E.
@@ 5.2.2 Analysis Logic
We investigated the Situational Dependence Bias in LLMs by analyzing their risk preferences across different scenarios . We examined choices in gain-framed scenarios to detect any situational bias in risk preferences. In the context of Loss Aversion, we scrutinized LLM responses in loss-framed scenarios to assess tendencies towards risk-aversion or risk-loving behaviors. To explore the Framing Effect, we translated scenarios from Chinese to English and monitored shifts in LLM preferences , observing how linguistic framing affects decisions. This approach aims to reveal if LLM outputs are biased by the linguistic construction of scenarios and choices.
5.2.3 Result
@@
The exploration of Risk-preference Bias entails the examination of LLM decisions across varied scenarios. The compiled results, particularly for gain-framed queries, are tabulated in Table 11. A predominant trend among most models is the exhibition of distinct risk preferences in disparate scenarios, indicative of a pronounced Situational Dependence Bias. Nonetheless, prefacing prompts with an instruction of the model’s risk-averse nature significantly attenuates this bias. In the context of loss-framed queries, some models exhibit a pronounced Loss Aversion Bias like GPT-4, as shown in LABEL:tab:Loss_Aversion_Bias. Moreover, refer to LABEL:tab:translation_difference, the translation of all queries into English precipitated notable discrepancies between the models’ responses to Chinese and English versions, underscoring a pronounced Framing Effect.
In particular, we have selected several representative cases for analysis, as shown in Figure 5. For Xuanyuan-13B, inducing Risk-Aversion does not alter its original preference distribution, but it exhibits a stronger Framing Effect. For GLM4, it performs well in terms of Framing Effect bias and can effectively switch preferences based on instructions. For Qwen-14B, it is capable of some preference distribution shift according to instructions, but also exhibits a significant Framing Effect.
6 Bias Tracker
After detecting Belief Bias and Risk-Preference Bias in LLMs, it was evident that Belief Bias had a more significant impact compared to Risk-Preference Bias. Risk-Preference Bias manifests primarily as the model’s inherent decision-making tendency, whereas Belief Bias results from the model’s misinterpretation of information, leading us to focus on analyzing Belief Bias formation.
To further investigate, we utilized Chain-of-Thought (COT) methods Wei et al. (2022) to engage the model’s System 2 thinking, aiming to enhance its in-depth analysis capabilities. We also scrutinized the model’s output, particularly the attention importance distribution across input tokens, enabling precise diagnosis of bias causes and informing subsequent bias correction strategies.
6.1 Slow Thinking
6.1.1 Methodology
To investigate the roots of score instability, whether due to inadequate reasoning or compromised rationality, we first employ a slow-thinking approach similar to CoT, prompting the model to generate reasoning before providing scores. This approach helps us study the causes of Belief Bias by enabling the model to produce reasoning for its evaluations, followed by the actual scoring. We utilize Bertopic Grootendorst (2022) for topic word extraction within the reasoning texts provided by the models. By clustering reasoning texts and extracting keywords, we analyze score discrepancies across different clusters, denoted by , to identify if certain thematic focuses lead to inconsistent evaluations. This detailed analysis helps us pinpoint specific causes of bias, thus informing future model optimizations and bias corrections.
6.1.2 Result
The application of the COT methodology, aimed at exploring the causes of financial bias in the model, yielded results as shown in H.2. The clustering and keyword analysis indicate that this intentional cognitive strategy improves the logical consistency and financial acumen of the model outputs. Integrating the provided thoughts into the previous sentence, the English expression would be: A comparison of the reasoning keywords for the top-performing GLM-4 model and the underperforming Baichuan2-7B model, illustrated in Figure 39 and Figure 39 respectively, reveals that the GLM-4 exhibits stronger logical coherence with primary keywords such as "termination", "transaction restructuring" and "withdrawal". In contrast, the Baichuan2-7B model’s logic is weaker, with primary keywords including "change", "decision", and "information". Moreover, the clustering of the models’ inferential outputs reveals substantial variations in the ratings assigned to each category. This divergence underscores that the irrationality exhibited by models in financial contexts stems more from their inherent cognitive processes than from their computational capabilities.
6.2 Attention Importance
6.2.1 Methodology
To verify whether the bias in the LLMs arises from an excessive focus on certain input tokens due to the training corpus, we examine the attention importance of LLMs for input sequences. Inspired by Wu et al. (2023), we define the importance of input token to output token as:
where is the context to generate by concatenating the prompt and the first tokens of response . omits token from , and is the conditional probability computed by the language model . We accelerate it with the first-order approximation:
where is the input word embedding of token . This approach helps us determine whether the LLM excessively focuses on specific financial entities, thereby more accurately diagnosing the specific causes of bias.
6.2.2 Result
Due to the limitations of Chinese tokens, we chose to use LLMs with BPE tokenizers, focusing on the well-performing MiniCPM-2B and the more biased Baichuan2-7B models, results are shown in Figure 6. We analyzed the attention each model pays to each input token in their outputs and found that Baichuan2-7B tends to focus more on financial company entities, industries, and their surrounding tokens. This excessive attention to irrelevant information further exacerbates the generation of financial biases.
7 Bias Antidote
7.1 Methodology
To eliminate Belief Bias in LLMs while preserving their original general capabilities, we employed four prompt-based methods for bias mitigation. 1) we utilized a CoT approach to enable the model to engage in slow thinking, thereby producing scores based on logical reasoning. 2) We implemented the S2A method Weston and Sukhbaatar (2023) to shield the model from irrelevant context, allowing for secondary reasoning before scoring. 3) We used Entity Replace to stabilize the model’s input. 4) We applied Fincausal relationship understanding based on knowledge to remove biases. For the last method, we extracted 200,000 pieces of financial causal knowledge about industries and individual stocks from past research reports and used a retrieval-augmented generation (RAG) approach to recall relevant causal information.
| Method | GLM-4 | ChatGLM3-Turbo | MiniCPM-2B | Xuanyuan2-6B | Baichuan2-7B |
| Direct | 0.598 | 1.067 | 1.409 | 13.999 | 28.106 |
[…]
7.2 Result
For Belief Bias, we employed four prompt-based elimination methods and selected five representative LLMs for bias elimination experiments, as shown in Table 1. For the CoT method, it performed well on the originally more biased models, as this reasoning approach enhances logical consistency in responses, thereby improving robustness and reducing bias. However, it performed poorly on models that were originally well-performing, as the increased output length due to the auto-regressive nature of these models resulted in amplified bias. For the S2A method, the effectiveness of increasing model output to reduce irrelevant attention depends on the model’s original capability; weaker models tend to exhibit greater bias. The Entity Replace method showed superior performance due to the substitution of financial topics. For the FinCausal method, each test data recalled four related causal knowledge entries, further enhancing the model’s reasoning ability to mitigate bias.
@@ 8 Discussion
Our study using the FBI framework aids in identifying and reducing irrationalities in finance sector models, enhancing understanding of LLM biases and logic, and applying methods to mitigate financial biases.
@@ 8.1 Model Size
The analysis in Section 5.1.3 reveals that within a specific model family, the degree of bias tends to decrease with the enlargement of model parameters, in line with the scaling law Kaplan et al. (2020). However, this trend is not consistent across different model family, where bias levels are also shaped by factors such as model design and training approaches.
@@ 8.2 Training Data
The FBI framework assessed financial rationality in general and financial LLMs. Financially-trained models, as seen in Figure 4, may exhibit higher score variability and risk inclination, potentially increasing financial irrationality. Models like ChatGLM2-6B and Qwen-7B show opposite industry biases (Figure 15-Figure 32), suggesting temporal biases in training data align with industry cycle rotation. Financial data, including potentially embellished research reports, can enhance LLMs for finance-specific NLP tasks but may also embed financial irrationality. Using these models in financial quantification could lead to unintended disruptions, demanding immediate attention.
@@ 8.3 Input Forms
The FBI framework utilized four prompt-based methods to eliminate biases, each of which produced certain effects, with the elimination being more pronounced when the model’s initial bias was more severe. As observed in Table 1, methods that increase output length, such as COT and S2A, can potentially exacerbate biases. The Entity Replace method, which reduces input, often yields better results, and the FinCausal method, which increases input, can further enhance the model’s ability for financial causal reasoning, achieving the effect of bias mitigation.
@@ 9 Conclusion
Our research introduces the FBI framework, a novel method for evaluating the financial rationality of LLMs in the intricate field of financial analysis. We rigorously examined 23 leading LLMs and revealed substantial differences in their financial rationality. By defining, detecting, analyzing, and mitigating financial biases, our study validates the capabilities and limitations of LLMs in financial contexts, offering reliable insights for their application in the finance sector. The advancement of LLMs towards greater financial acumen and reduced bias is essential for their dependable use in financial analysis, paving the way for a future where AI-generated insights are confidently and precisely applied in the financial industry.
Limitation
Our financial rational analysis focuses on biases towards the awareness of the Chinese A-share market, which may vary depending on culture and region, resulting in research findings that may not be generalizable to other situations. Meanwhile, we introduced other models during model evaluation, which may lead to the introduction of other biases.
Ethics Statement
This paper honors the EMNLP Code of Ethics. The dataset used in the paper does not contain any private information. All annotators have received enough labor fees corresponding to their amount of annotated instances. The code and data are open-sourced under the MIT license.
@@ Appendix A Cognitive Bias
| Belief Bias | Limited Attention | The brain has two systems when working: fast thinking and slow thinking. It uses intuition to deal with things quickly. |
| Representativeness bias | When making probability estimates, people tend to focus on certain representative features, ignoring environmental probabilities and sample size. | |
| Anchoring effect | Decision-making is often influenced by the first information received, like an anchor sinking to the bottom of the sea. | |
| Overconfidence | Belief that one’s knowledge is more accurate than the facts; one’s information is given more weight. | |
| Risk-Preference Bias | Situational dependence bias | The effect of a stimulus depends largely on the context in which it occurs. |
| Loss aversion | Sensitivity to losses exceeds gains of equal value. | |
| Framing effect | Different descriptions of an objectively identical problem lead to different decision-making judgments. |
Appendix B Company Profile
Appendix C Event Type
We have sorted out the types of events that can affect a company’s stock price based on the regular patterns of the Chinese A-Share stock market, and finally sorted out four categories, totaling 16 types of events. The detailed content is shown in Table 3.
@@ Appendix C Event Type
| Corporate Governance and Equity Changes | Major Asset Restructuring | The process of recombining, adjusting, and allocating the distribution status of enterprise assets among the owners, controllers, and external economic entities. |
| Equity Incentive | By conditionally granting employees partial shareholder rights, a sense of ownership is fostered, forming a community of interests with the company. | |
| Increase or Decrease in Shareholder Holdings | Changes in the shareholder holdings of company stocks. | |
| Buy-back | The act of a listed company using cash or other means to repurchase its shares from the stock market. | |
| Circulation of Restricted Stock | Restricted shares become freely tradable in the secondary market after the commitment period. | |
| Financial Reports and Earnings Expectations | Performance Report | Regular preparation by each responsibility center to evaluate and assess performance, serving as the basis for future budget preparation. |
| Market Behavior and Announcements | Private Placement | Targeted issuance of bonds or stocks to a select group of senior institutional or individual investors. |
| Transfer of Shares | Listed companies transfer their provident fund to share capital in proportion or issue bonus shares accordingly. | |
| Stock Price Fluctuations | Sudden large inflows and outflows of funds lead to increased volatility in stock prices. | |
| Business Dynamics | Updates on enterprises and their surroundings, using major production and sales information to promote corporate brand and image. | |
| Negative Events and Risk Management | Dispute | Disputes between companies or between companies and individuals. |
| Investigation | Filing an investigation signifies a basic determination of illegal facts, allowing for compulsory measures and official initiation of investigation procedures. | |
| Violation Penalties | Punishments for enterprises violating regulations of regulatory bodies. | |
| Litigation and Arbitration | Litigation and arbitration for contract disputes and other property rights disputes between enterprises. | |
| Security | Enterprises providing guarantees for loans and other matters for other enterprises. |
@@ Appendix D Models
We have selected a total of 23 financial and general LLMs oriented by Chinese and English, with specific details shown in LABEL:tab:models.
@@
| MiniCPM-2B | True | 2B | False | local |
| Baichuan-13B | True | 13B | False | local |
| DISC-FinLLM | True | 13B | True | local |
| Baichuan2-7B | True | 7B | False | local |
| Baichuan2-13B | True | 13B | False | local |
| ChatGLM2-6B | True | 6B | False | local |
| ChatGLM3-6B | True | 6B | False | local |
| ChatGLM3-Turbo | True | 33B | False | API |
| GLM-4 | True | Unknown | False | API |
| InternLM2-7B | True | 7B | False | local |
| InternLM2-20B | True | 20B | False | local |
| LLaMA2-7B | False | 7B | False | local |
| LLaMA2-13B | False | 13B | False | local |
| Qwen-7B | True | 7B | False | local |
| Qwen-14B | True | 14B | False | local |
| FinQwen | True | 14B | True | local |
| Qwen-72B | True | 72B | False | local |
| Qwen-max | True | 72B | False | API |
| Xuanyuan-13B | True | 13B | True | local |
| Xuanyuan-70B | True | 70B | True | local |
| Xuanyuan2-6B | True | 6B | True | local |
| GPT-3.5 | False | Unknown | False | API |
| GPT-4 | False | Unknown | False | API |